Inductor L = 100mH , resistance =100 W AND battery = 100 V are connected in series. After long time battery is disconnected current in it after 10-3 s?? ….….1/e A, A e, 1 A

Grade:

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The current at any instant of time is given by:
I\left ( t \right )=\frac{V}{R}\left ( 1-e^{\frac{-Rt}{L}} \right )
hence, the current would be given by:
I\left ( t \right )=\frac{V}{R}\left ( 1-e^{\frac{-Rt}{L}} \right )=\frac{100}{100}\left ( 1-e^{\frac{-100\times 10^{-3}}{100\times 10^{-3}}} \right )=\left ( 1-e^{-1} \right ) amp
Regards
Sumit
Saurabh Kumar
9 years ago
For the discharging, through R, put the value and find the answer.